OpenStudy (anonymous):

y=3x^2+6x-10 find the value of x where y has the smallest value. What is the smallest value of y?

OpenStudy (anonymous):

look for vertex

OpenStudy (anonymous):

notice the parabola is opening up and so its smallest y value is the vertex. We now it opens up because the # in front of x^2 is positive

OpenStudy (anonymous):



OpenStudy (anonymous):

the x-value for the vertex is x=-b/2a where we have an equation of the form \[ax^2+bx+c=f(x) \]

OpenStudy (anonymous):

identify b and a

OpenStudy (anonymous):

but how do I get rid of the 3 to make x^2 by itself?

OpenStudy (anonymous):

you don't need to do any separating. to find the vertex it is all plug and play.

OpenStudy (anonymous):